Instructions

Please fill the user’s information on the request form, and sign it at the bottom. The user’s supervisor must fill the supervisor’s information and sign the form. In case the user holds a permanent position within the Institution specified on the form, he/she may fill and sign as his/her own supervisor.

Please return the completed form by post to

Alfredo Ferrari EN/STI, Mailbox Z04500 CERN CH-1211 Geneva 23 Switzerland fax: +41-227669474

Sending of a copy by fax may speed up the procedure, however the original form must also be sent by standard mail.

Request form for the FLUKA code

April 28, 2010

Single-User , Single-site information

Name  Nicholas Bolibruch
FLUKA user id fuid-2139

Institution

Address

e-mail  nicholas.bolibruch@lightsource.ca

Liason Officer or Supervisor

Name

e-mail

By signature below, the subscriber acknowledges and agrees to comply with the FLUKA User license and the associated Author Requests as reported below.

The subscriber acknowledges and agrees that the FLUKA software is licensed, on a non-exclusive basis, for use only by the identified single user only on computers owned or leased by the user’s Institution as designated above. The license expires automatically when the user leaves his/her position in the specified Institution. Access by anyone other than the identified user is strictly prohibited. The subscriber declares and agrees that the FLUKA code will not be used directly or indirectly in any or more of the following areas

• Weapons or their precursors, and associated hardware, software or technical data

• Commercial activities

• Training, unless fee-less and restricted to academic/educational audience.

The licensee declares that he/she has carefully read the release notes and he/she accepts the associated recommendations. In particular the licensee acknowledges that

• The availability of the source code shall not be exploited for tentative builds on other architectures or with different compilers/compiler options than the ones recommended by the development team.

• The licensee shall regularly update his FLUKA version according to the FLUKA release notes. The availability of the source code shall not be exploited to circumvent the expiration date.

FLUKA User license:

By downloading and/or using FLUKA, you agree with the following license/conditions/requests:

Copyright statement and license conditions

Copyright Italian National Institute for Nuclear Physics (INFN) and European Organization for Nuclear Research (CERN) (”the FLUKA copyright holders”), 1989-2008.

All rights not expressly granted under this license are reserved.

This software results from work performed by Alberto Fasso‘, Alfredo Ferrari, Johannes Ranft and Paola Sala.

INFN and CERN are the exclusive source of distribution of the code, bug fixes and documentation of the FLUKA software. Each official version of FLUKA is identified by a numbering scheme specifying major and minor releases.

The FLUKA Coordination Committee or its delegates are able to grant any of the permissions noted in this License Agreement as requiring a specific consent. Any such consent may only be granted in writing.

Installation, use, reproduction, display of the FLUKA software (”FLUKA”), in source and binary forms, are permitted free of charge on a non-exclusive basis for internal scientific, non-commercial and non-weapon-related use by non-profit organizations only. Any exercise of these rights is subject to the following conditions:

1. Insertion of the FLUKA code, in whole or in part, into other codes, or its translation into any other computer language are possible only after obtaining prior written permission. Modifications of the FLUKA code are permitted for use by the licensee only, unless authorized in written.

2. FLUKA is non-transferable, non-sub-licensable and may not be distributed in any way, without express written consent, whether in original or modified form. Sitewise or collaboration-wise conditions can be agreed with the FLUKA Coordination Committee.

3. Notwithstanding the above, the licensee may modify and sub-license FLUKA User Routines to third parties in so far as their purpose is limited to the adaptation of input and output interfaces of FLUKA and their modification does not circumvent, replace, add to or modify any of the functions of FLUKA, or extract specific isolated results from any of the individual internal physics models embedded within FLUKA.

4. The licensee shall forthwith license all its modifications of FLUKA to the FLUKA copyright holders, at no cost and with no limitation of use. The licensee acknowledges that the FLUKA copyright holders may insert such modifications into future releases of FLUKA, subject to appropriate acknowledgment of the licensee’s contribution.

5. Any publication by the licensee with respect to FLUKA or results obtained from it shall explicitly acknowledge FLUKA by quoting its set of references and the FLUKA copyright holders. The licensee shall not without prior written permission publish documents or results based on a modified form of FLUKA, unless the modification exclusively concerns User Routines for the adaptation of its input and output interfaces which comply with the same restrictions, as defined in section 3) as those which apply to sub-licensing. Any publication of documents or results shall be based only on official FLUKA versions as obtained from the FLUKA website (http://www.fluka.org) or from any authorized mirror. Publication here implies any legal publication to any third party, whether verbal, electronic, visual, in writing or otherwise.

6. The licensee shall ensure that the FLUKA references, copyright statement and license conditions are not altered or removed from FLUKA. Any integration of any portion of FLUKA, in modified or in unmodified form, into any other software package must preserve the internal copyright notices in those portions of FLUKA that have been employed, and must reproduce such notices within any additional global notices included along or embedded within the software into which FLUKA has been integrated. Any portion of FLUKA so integrated, whether modified or unmodified shall continue to be subject to these license conditions.

7. Nothing in this license shall be construed as to grant any rights in any of the FLUKA versions since 1989. In addition, users are not permitted to circumvent any protection in prior distributions of FLUKA that provided for a preset expiration date of the code

8. Versions or parts of the FLUKA source code, entrusted to individuals or groups prior to the enactment of the CERN-INFN Collaboration Agreement, which are listed in Chapter 5 of Annex 1 of the EP-AB-INFN Scientific Agreement (19-02-2003), together with the agreed conditions of use, are subject to this License Agreement in addition to any other restrictions on the scope of use that may have been part of the initial use grant.

9. Commercial use of FLUKA, outside the scope of this license, must be negotiated with the copyright holders

10. DISCLAIMER

THIS SOFTWARE IS PROVIDED BY THE FLUKA COPYRIGHT HOLDERS ”AS IS” AND ANY EXPRESS OR IMPLIED WARRANTIES, INCLUDING, BUT NOT LIMITED TO, IMPLIED WARRANTIES OF MERCHANTABILITY, OF SATISFACTORY QUALITY, AND FITNESS FOR A PARTICULAR PURPOSE OR USE ARE DISCLAIMED. THE FLUKA COPYRIGHT HOLDERS AND THE AUTHORS MAKE NO REPRESENTATION THAT THE SOFTWARE AND MOD-IFICATIONS THEREOF, WILL NOT INFRINGE ANY PATENT, COPYRIGHT, TRADE SECRET OR OTHER PROPRIETARY RIGHT.

11. LIMITATION OF LIABILITY

THE FLUKA COPYRIGHT HOLDERS AND THE AUTHORS SHALL HAVE NO LIABILITY FOR DIRECT, INDIRECT, SPECIAL, INCIDENTAL, CONSE-QUENTIAL, EXEMPLARY, OR PUNITIVE DAMAGES OF ANY CHARACTER INCLUDING, WITHOUT LIMITATION, PROCUREMENT OF SUBSTITUTE GOODS OR SERVICES, LOSS OF USE, DATA OR PROFITS, OR BUSINESS INTERRUPTION, HOWEVER CAUSED AND ON ANY THEORY OF CON-TRACT, WARRANTY, TORT (INCLUDING NEGLIGENCE), PRODUCT LIA-BILITY OR OTHERWISE, ARISING IN ANY WAY OUT OF THE USE OF THIS SOFTWARE, EVEN IF ADVISED OF THE POSSIBILITY OF SUCH DAMAGES.

REQUESTS MADE BY THE FLUKA AUTHORS TO ALL USERS

• All licensees are requested to report as soon as practical to the Authors any errors or bugs found in any portion of FLUKA and its accompanying documentation.

• All licensees are requested to forward copies of all comparisons that they may make between FLUKA results and data or other codes as soon as practical. The Authors agree to keep any such communications confidential unless otherwise notified by the contributing user.

• The Authors reserve the right of publishing any benchmarking and/or comparisons of the distinct separate performance of the individual internal models that are embedded into FLUKA, whether the comparisons are with data or with other codes. The Authors would also like to convey a general willingness to conduct any such benchmarking efforts either upon request or in collaboration with interested parties. In case of doubt please contact the Authors.

• Users should exclusively download FLUKA from the official FLUKA website (http://www.fluka.org) or one of the authorized mirror sites. Users are invited to regularly update their FLUKA version to profit for improvements and bug fixes.

• Users are invited to use reasonably updated versions of the code for publications.

Publications of results based on those FLUKA versions that are declared unsupported and obsolete on the official FLUKA website shall be avoided.

• Users should address any request of consent to one member of the FLUKA Coordinating Committee, which at present is composed as follows:

Giuseppe Battistoni Giuseppe.Battistoni@mi.infn.it (chairman) Michael Doser Michael.Doser@cern.ch Roberto Losito Roberto.Losito@cern.ch Johannes Ranft Johannes.Ranft@cern.ch Paola Sala Paola.Sala@mi.infn.it

In accordance with the User License, the use of the FLUKA code shall be acknowledged explicitly by quoting the following and only the following set of references:

• A. Fassò, A. Ferrari, J. Ranft, P.R. Sala, “ FLUKA a multi-particle transport code ”, CERN 2005-10 (2005), INFN/TC 05/11, SLAC-R-773

• G. Battistoni, S. Muraro, P.R. Sala, F. Cerutti, A. Ferrari, S. Roesler, A. Fasso‘, J. Ranft, ”The FLUKA code: Description and benchmarking”, Proceedings of the Hadronic Shower Simulation Workshop 2006, Fermilab 6–8 September 2006, M. Albrow, R. Raja eds., AIP Conference Proceeding 896, 31-49, (2007)

Signature:

print name  signature  date

and as supervisor:

print name  signature  date